# Debate or Vote: Which Yields Better Decisions in Multi-Agent Large Language Models?

### v1

@@ Acknowledgement
The authors would like to thank Leitian Tao and Xuanming Zhang for their valuable comments on the manuscript. Hyeong Kyu Choi and Yixuan Li are supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon. Xiaojin Zhu was supported in part by NSF grants 2202457, 2331669, 1836978, 2023239, ARO MURI W911NF2110317, and AF CoE FA9550-18-1-0166.
@@ B.3 Persona Prompts
Appendix C Proofs
C.1 Proof of Theorem 1
@@ C.2 Proof of Lemma 1
C.3 Proof of Theorem 2
Theorem 2. (Martingale Behavior of Multi-Agent Debate) For agent $i$, let $p_{t}:=\bar{{\theta}}_{i,t}^{(1)}$ denote its belief in the correct answer at debate round $t$. Under Bayesian conjugacy,
$p_{t}:=\bar{\theta}_{i,t}^{(1)}=\frac{\alpha_{i,t}^{(1)}}{\sum_{j=1}^{K}\alpha_{i,t}^{(j)}}=\frac{\alpha_{i,t-1}^{(1)}+c_{i,t}^{(1)}}{\sum_{j=1}^{K}(\alpha_{i,t-1}^{(j)}+c_{i,t}^{(j)})}.$
@@ C.3 Proof of Theorem 2
Then, sequence $\{p_{t}\}_{t\geq 0}$ forms a martingale. That is, the expected belief at the next round equals the current belief:
$\mathbb{E}[p_{t}\mid p_{t-1},\ldots,p_{0}]=p_{t-1},\;\forall_{t\geq 0}.$
@@
Proof of Theorem 2. We prove that the sequence $\{p_{t}\}_{t\geq 0}$, where $p_{t}=\bar{\theta}_{i,t}^{(1)}$, is a martingale by showing that the conditional expectation of the belief in the correct answer at round $t+1$ given the filtration up to round $t$ equals the current belief. A sequence $\{X_{t}\}_{t\geq 0}$ is a martingale if $\mathbb{E}[X_{t+1}\mid\mathcal{F}_{t}]=X_{t}$, where $\mathcal{F}_{t}=\sigma(\bar{\theta}_{i,0}^{(1)},\ldots,\bar{\theta}_{i,t}^{(1)})$ represents the history of beliefs.
By the problem statement, agent $i$ initializes its belief with $\boldsymbol{\theta}_{i,0}\sim\mathrm{Dirichlet}(\boldsymbol{\alpha}_{i,0})$, where $\boldsymbol{\alpha}_{i,0}=(\alpha_{1},\ldots,\alpha_{K})$ with $\alpha_{j}>0$. The initial expected belief is:
$p_{0}=\bar{\theta}_{i,0}^{(1)}=\frac{\alpha_{i,0}^{(1)}}{\sum_{j=1}^{K}\alpha_{i,0}^{(j)}}.$
At each round $t>0$, agent $i$ observes the responses $\{y_{j,t-1}\mid j\in\mathcal{N}(i)\}$ from its neighbors, aggregates them into a count vector $\mathbf{c}_{i,t}=(c_{i,t}^{(1)},\ldots,c_{i,t}^{(K)})$, where $c_{i,t}^{(m)}=\sum_{j\in\mathcal{N}(i)}\mathbf{1}[y_{j,t-1}=m]$, and updates its parameter vector as:
$\boldsymbol{\alpha}_{i,t}=\boldsymbol{\alpha}_{i,t-1}+\mathbf{c}_{i,t}.$
The belief in the correct answer (assumed to be class 1) at round $t$ is a random variable $\bar{\theta}_{i,t}^{(1)}$, and its expectation is:
$p_{t}=\bar{\theta}_{i,t}^{(1)}=\frac{\alpha_{i,t}^{(1)}}{\sum_{j=1}^{K}\alpha_{i,t}^{(j)}}.$
To verify the martingale property, we compute the conditional expectation of $\bar{\theta}_{i,t+1}^{(1)}$ given $\mathcal{F}_{t}$. At round $t+1$, agent $i$ observes $\{y_{j,t}\mid j\in\mathcal{N}(i)\}$, forming a new count vector $\mathbf{c}_{i,t+1}$, where $c_{i,t+1}^{(1)}=\sum_{j\in\mathcal{N}(i)}\mathbf{1}[y_{j,t}=1]$. The updated parameter is:
$\boldsymbol{\alpha}_{i,t+1}=\boldsymbol{\alpha}_{i,t}+\mathbf{c}_{i,t+1},$
and the belief is:
$\bar{\theta}_{i,t+1}^{(1)}=\frac{\alpha_{i,t}^{(1)}+c_{i,t+1}^{(1)}}{\sum_{j=1}^{K}(\alpha_{i,t}^{(j)}+c_{i,t+1}^{(j)})}.$
The conditional expectation is:
$\mathbb{E}[\bar{\theta}_{i,t+1}^{(1)}\mid\mathcal{F}_{t}]=\mathbb{E}\left[\frac{\alpha_{i,t}^{(1)}+c_{i,t+1}^{(1)}}{\sum_{j=1}^{K}(\alpha_{i,t}^{(j)}+c_{i,t+1}^{(j)})}\mid\mathcal{F}_{t}\right].$
Since $\boldsymbol{\alpha}_{i,t}$ is deterministic given $\mathcal{F}_{t}$, and $c_{i,t+1}^{(j)}$ depends on the random responses $y_{j,t}$, we need the expectation of the counts. Each $y_{j,t}\sim\text{Categorical}(\bar{\theta}_{j,t})$, and $\bar{\theta}_{j,t}^{(1)}$ is the belief of neighbor $j$ at round $t$, which is in $\mathcal{F}_{t}$ (assuming $\bar{\theta}_{j,t}$ is updated and observable up to $t$). Thus:
$\mathbb{E}[\mathbf{1}[y_{j,t}=1]\mid\mathcal{F}_{t}]=\bar{\theta}_{j,t}^{(1)},$
and:
$\mathbb{E}[c_{i,t+1}^{(1)}\mid\mathcal{F}_{t}]=\mathbb{E}\left[\sum_{j\in\mathcal{N}(i)}\mathbf{1}[y_{j,t}=1]\mid\mathcal{F}_{t}\right]=\sum_{j\in\mathcal{N}(i)}\bar{\theta}_{j,t}^{(1)},$
$\mathbb{E}[c_{i,t+1}^{(j)}\mid\mathcal{F}_{t}]=\sum_{j\in\mathcal{N}(i)}\bar{\theta}_{j,t}^{(j)},$
and the denominator’s expected value is:
$\mathbb{E}\left[\sum_{j=1}^{K}(\alpha_{i,t}^{(j)}+c_{i,t+1}^{(j)})\mid\mathcal{F}_{t}\right]=\sum_{j=1}^{K}\alpha_{i,t}^{(j)}+\sum_{j=1}^{K}\sum_{l\in\mathcal{N}(i)}\bar{\theta}_{l,t}^{(j)}.$
However, since $\bar{\theta}_{j,t}$ is a probability vector summing to 1, and $c_{i,t+1}^{(j)}$ counts over $\mathcal{N}(i)$, the total expected count is $|\mathcal{N}(i)|$. The correct approach is to recognize that:
$\bar{\theta}_{i,t+1}^{(1)}=\frac{\alpha_{i,t}^{(1)}+c_{i,t+1}^{(1)}}{\alpha_{i,t}^{0}+|\mathcal{N}(i)|},$
where $\alpha_{i,t}^{0}=\sum_{j=1}^{K}\alpha_{i,t}^{(j)}$. The conditional expectation becomes:
$\mathbb{E}[\bar{\theta}_{i,t+1}^{(1)}\mid\mathcal{F}_{t}]=\frac{\alpha_{i,t}^{(1)}+\mathbb{E}[c_{i,t+1}^{(1)}\mid\mathcal{F}_{t}]}{\alpha_{i,t}^{0}+|\mathcal{N}(i)|}.$
Substituting the expected count:
$\mathbb{E}[c_{i,t+1}^{(1)}\mid\mathcal{F}_{t}]=\sum_{j\in\mathcal{N}(i)}\bar{\theta}_{j,t}^{(1)},$
so:
$\mathbb{E}[\bar{\theta}_{i,t+1}^{(1)}\mid\mathcal{F}_{t}]=\frac{\alpha_{i,t}^{(1)}+\sum_{j\in\mathcal{N}(i)}\bar{\theta}_{j,t}^{(1)}}{\alpha_{i,t}^{0}+|\mathcal{N}(i)|}.$
For this to equal $\bar{\theta}_{i,t}^{(1)}=\frac{\alpha_{i,t}^{(1)}}{\alpha_{i,t}^{0}}$, we need:
$\frac{\alpha_{i,t}^{(1)}+\sum_{j\in\mathcal{N}(i)}\bar{\theta}_{j,t}^{(1)}}{\alpha_{i,t}^{0}+|\mathcal{N}(i)|}=\frac{\alpha_{i,t}^{(1)}}{\alpha_{i,t}^{0}},$
which implies:
$\alpha_{i,t}^{(1)}+\sum_{j\in\mathcal{N}(i)}\bar{\theta}_{j,t}^{(1)}=\frac{\alpha_{i,t}^{(1)}(\alpha_{i,t}^{0}+|\mathcal{N}(i)|)}{\alpha_{i,t}^{0}},$
$\sum_{j\in\mathcal{N}(i)}\bar{\theta}_{j,t}^{(1)}=\alpha_{i,t}^{(1)}\cdot\frac{|\mathcal{N}(i)|}{\alpha_{i,t}^{0}},$
$\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}\bar{\theta}_{j,t}^{(1)}=\frac{\alpha_{i,t}^{(1)}}{\alpha_{i,t}^{0}}=\bar{\theta}_{i,t}^{(1)}.$
@@
Appendix D Special Case of Theorem 1
Rewrite in terms of $\bar{X}$:
$P(\bar{X}\leq p-\epsilon)\leq\exp(-2n\epsilon^{2})$
We want the majority vote to be successful, i.e., $\bar{X}>\frac{1}{2}$.
First, consider the complementary event:
and set $\epsilon=p-\frac{1}{2}$:
$p-\epsilon=\frac{1}{2}\implies\epsilon=p-\frac{1}{2}$
Since $p>\frac{1}{2}$, $\epsilon>0$. Substitute $\epsilon$:
Thus, the probability of a successful majority vote is:
@@ Appendix E Martingale Process Empirical Investigation
Table 6: Raw mean accuracy values for Figure 4. Top row is from Qwen2.5-7B-Instruct, and the bottom row is from Llama3.1-8B-Instruct.
@@ Appendix G Closed-source LLM Evaluation
Appendix H Limitations and Future Works

### v2

@@ Acknowledgement
The authors would like to thank Leitian Tao and Xuanming Zhang for their valuable comments on the manuscript. Hyeong Kyu Choi and Sharon Li are supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon. Xiaojin Zhu was supported in part by NSF grants 2202457, 2331669, 1836978, 2023239, ARO MURI W911NF2110317, and AF CoE FA9550-18-1-0166.
@@ B.3 Persona Prompts
Appendix C Special Case of Theorem 1
Rewrite in terms of $\bar{X}$:
$P(\bar{X}\leq p-\epsilon)\leq\exp(-2n\epsilon^{2})$
We want the majority vote to be successful, i.e., $\bar{X}>\frac{1}{2}$.
First, consider the complementary event:
and set $\epsilon=p-\frac{1}{2}$:
$p-\epsilon=\frac{1}{2}\implies\epsilon=p-\frac{1}{2}$
Since $p>\frac{1}{2}$, $\epsilon>0$. Substitute $\epsilon$:
Thus, the probability of a successful majority vote is:
[…]
Appendix D Proofs
D.1 Proof of Theorem 1
@@ D.2 Proof of Lemma 1
D.3 Proof of Theorem 2
Theorem 2. (Martingale Behavior of Multi-Agent Debate)
For any agent $i$ at round $t>0$, if
$\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}p_{j,t-1}=p_{i,t-1},$
@@ D.3 Proof of Theorem 2
then sequence $\{p_{i,t}\}_{t\geq 0}$ forms a martingale. That is, the expected belief at the next round equals the current belief:
$\mathbb{E}[p_{i,t}\mid\boldsymbol{\alpha}_{t-1}]=p_{i,t-1}.$
Proof of Theorem 2. For agent $i$ with neighbors $\mathcal{N}(i)$, let $p_{i,t}:=\bar{{\theta}}_{i,t}^{(1)}$ denote its belief in the correct answer at debate round $t$. Under Bayesian conjugacy,
$p_{i,t}:=\bar{\theta}_{i,t}^{(1)}=\frac{\alpha_{i,t}^{(1)}}{\sum_{j=1}^{K}\alpha_{i,t}^{(j)}}=\frac{\alpha_{i,t-1}^{(1)}+c_{i,t}^{(1)}}{\sum_{j=1}^{K}(\alpha_{i,t-1}^{(j)}+c_{i,t}^{(j)})}.$
@@
Here, we prove that the sequence $\{p_{i,t}\}_{t\geq 0}$ is a martingale by showing that the conditional expectation of the belief in the correct answer at round $t$ given the filtration up to round t1t-1 equals the current belief. A sequence $\{X_{t}\}_{t\geq 0}$ is a martingale if $\mathbb{E}[X_{t}\mid\mathcal{F}_{t-1}]=X_{t-1}$, where $\mathcal{F}_{t-1}=(\boldsymbol{\alpha}_{0},\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{t-1})$ represents the history of belief parameters.
By definition, we can write an agent’s belief in the correct answer at timestep t1t-1 and $t$ as:
$p_{i,t-1}$ $=\frac{\alpha_{i,t-1}^{(1)}}{\sum_{j=1}^{K}\alpha_{i,t-1}^{(j)}}$ (2)
$p_{i,t}$ $=\frac{\alpha_{i,t-1}^{(1)}+c_{i,t}^{(1)}}{\sum_{j=1}^{K}\alpha_{i,t-1}^{(j)}+|\mathcal{N}(i)|},$ (3)
where $c_{i,t}^{(1)}$ is the number of neighboring agents that has selected the correct answer at the previous debate round. Then, the expectation of $p_{t}$ conditioned on the filtration $\mathcal{F}_{t-1}$ is:
$\mathbb{E}[p_{i,t}\;|\;\mathcal{F}_{t-1}]$ $=\mathbb{E}\left[\frac{\alpha_{i,t-1}^{(1)}+c_{i,t}^{(1)}}{\sum_{j=1}^{K}\alpha_{i,t-1}^{(j)}+|\mathcal{N}(i)|}\;|\;\mathcal{F}_{t-1}\right]$ (4)
$=\frac{\alpha_{i,t-1}^{(1)}+\mathbb{E}[c_{i,t}^{(1)}\;|\;\mathcal{F}_{t-1}]}{\sum_{j=1}^{K}\alpha_{i,t-1}^{(j)}+|\mathcal{N}(i)|}$ (5)
$=\frac{\alpha_{i,t-1}^{(1)}+\mathbb{E}[\sum_{j\in\mathcal{N}(i)}\mathbf{1}\{y_{j,t-1}=1\}\;|\;\mathcal{F}_{t-1}]}{\sum_{j=1}^{K}\alpha_{i,t-1}^{(j)}+|\mathcal{N}(i)|}$ (6)
$=\frac{\alpha_{i,t-1}^{(1)}+\sum_{j\in\mathcal{N}(i)}p_{j,t-1}}{\sum_{j=1}^{K}\alpha_{i,t-1}^{(j)}+|\mathcal{N}(i)|}$ (7)
$=\frac{\alpha_{i,t-1}^{(1)}+|\mathcal{N}(i)|\,p_{i,t-1}}{\sum_{j=1}^{K}\alpha_{i,t-1}^{(j)}+|\mathcal{N}(i)|}\qquad\text{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(by the condition in Theorem 2)}.$ (8)
From equation (2), we have
$\alpha_{i,t-1}^{(1)}=p_{i,t-1}\sum_{j=1}^{K}\alpha_{i,t-1}^{(j)},$ (9)
and by plugging equation (9) into equation (8), we arrive at
$\mathbb{E}[p_{i,t}\;|\;\mathcal{F}_{t-1}]=p_{i,t-1}.$ (10)
@@
D.3.1 When does (1) in Theorem 2 Hold?
In order for a multi-agent debate to be a martingale process, the condition in Theorem 2, equation (1), which states the relation between the belief of an agent with its peers’ beliefs, should hold. Here, we expand our discussion to exploring debate topologies that satisfy (1), therefore the martingale property.
Proposition 1. (Sufficient Conditions for (1))
If the debate topology consists of agents that either (i) are homogeneous and fully-connected, or (ii) are fully-isolated, or (iii) form isolated cliques whose agents in each clique are homogeneous, then the debate process is a martingale.
proof of Proposition 1. We prove one by one that each case satisfies (1).
(i) Homogeneous agents with fully-connected communication topology
@@ D.3.1 When does (1) in Theorem 2 Hold?
(ii) Fully-isolated structure with $\mathcal{N}(i)=\{i\}$ for all $i$
[…]
(iii) Isolated cliques, where agents in a clique are homogeneous
[…]
Remark 2.
Proposition 1 provides several sufficient conditions under which equation (1) holds. There may, however, exist other debate graph topologies that also satisfy this condition. A complete characterization of such cases is left for future work.
D.3.2 When does (1) in Theorem 2 Not hold?
In addition to the cases where the condition in Theorem 2 hold, it is also important to examine the debate topologies under which (1) does not hold. Here, we discuss two cases: line graph and ring graph.
1. Line graph
A debate topology that consists of three agents connected as a line may not satisfy (1), as shown below:
123
[…]
In this case, $p_{2,1}\neq\frac{1}{3}\left(p_{1,1}+p_{2,1}+p_{3,1}\right)$. Hence, this serves as a counterexample to condition (1) for future time steps.
2. Ring graph
A debate topology that consists of four agents forming a ring graph may not satisfy (1), as shown below:
1234
[…]
In this case, $p_{3,1}\neq\frac{1}{3}\left(p_{2,1}+p_{3,1}+p_{4,1}\right)$. Hence, this serves as a counterexample to (1). Notably, this ring graph corresponds to the Sparse MAD baseline [10]. Although this topology does not theoretically guarantee the martingale property, we empirically demonstrate in Appendix E that it holds approximately. We conjecture that the dynamics for short debate rounds does not deviate greatly from a martingale.
D.4 Collective Intelligence
If the agents are heterogeneous and fully-connected, the debate is not necessarily a martingale since the condition (1) for individual agent beliefs may not hold. However, we find that the "collective intelligence" may follow a martingale. Let’s define the collective belief of the correct answer as:
$q_{t}:=\frac{\sum_{i}\alpha_{i,t}^{(1)}}{\sum_{i,k}\alpha_{i,t}^{(k)}}.$ (18)
Then, the expected collective belief is:
$\mathbb{E}[q_{t}\;|\;\mathcal{F}_{t}]$ $=\frac{\sum_{i}\left(\alpha_{i,t-1}^{(1)}+\mathbb{E}[c_{i,t}^{(1)}\;|\;\mathcal{F}_{t-1}]\right)}{\sum_{i,k}\alpha_{i,t-1}^{(k)}+\sum_{i}|\mathcal{N}(i)|}$ (19)
$=\frac{\sum_{i}\left(\alpha_{i,t-1}^{(1)}+\mathbb{E}[\sum_{j\in\mathcal{N}(i)}\mathbf{1}\{y_{j,t-1}=1\}\;|\;\mathcal{F}_{t-1}]\right)}{\sum_{i,k}\alpha_{i,t-1}^{(j)}+\sum_{i}|\mathcal{N}(i)|}.$ (20)
This collective belief will be a martingale if and only if
$\frac{\sum_{i}\sum_{j\in\mathcal{N}(i)}\mathbb{E}[\mathbf{1}\{y_{j,t-1}=1\}\;|\;\mathcal{F}_{t}]}{\sum_{i}|\mathcal{N}(i)|}=q_{t-1}$ (21)
is satisfied in (20). This condition is equivalent to
$\frac{\sum_{i}\sum_{j\in\mathcal{N}(i)}p_{j,t-1}}{\sum_{i}|\mathcal{N}(i)|}=\frac{\sum_{i}\alpha_{i,t-1}^{(1)}}{\sum_{i,k}\alpha_{i,t-1}^{(k)}}.$ (22)
Then, a sufficient condition for this to hold is, for all $i$,
$\begin{cases}\sum_{k}\alpha_{i,t-1}^{(k)}=\text{Constant},\\ |\mathcal{N}(i)|=n,\end{cases}$
meaning that the magnitude of the belief vector should be constant across agents, and that all agents should be connected to each other. When these conditions are met, $\{q_{t}\}_{t\geq 0}$ follows a martingale process.
@@ Appendix E Martingale Process Empirical Investigation
As further investigation, we directly evaluate the martingale process by comparing the $p_{i,t-1}$, as measured by the mean accuracy values across agents at step t1t-1, with the expected value of the next step $\mathbb{E}[p_{i,t}\;|\;\mathcal{F}_{t-1}]$, which is measured by running 5 independent runs on the subsequent debate round. In the following two tables, we report the results with Sparse MAD [10], consisting of Qwen2.5-7B agents and Llama3.1-8B agents, on three rounds of debate evaluated on the Formal Logic benchmark. We observe that $p_{i,t-1}\approx\mathbb{E}[p_{i,t}\;|\;\mathcal{F}_{t-1}]$ in both cases.
Table 6: Raw mean accuracy values for Figure 4. Top row is from Qwen2.5-7B-Instruct, and the bottom row is from Llama3.1-8B-Instruct. The MAD structure is the fully connected Decentralized MAD [2].
@@
Table 7: Qwen2.5-7B-Instruct on Formal Logic
  Round $t$   $p_{i,t-1}$   $\mathbb{E}[p_{i,t}\;|\;\mathcal{F}_{i,t-1}]$
  1   0.363   0.386 (0.024)
  2   0.400   0.410 (0.011)
  3   0.371   0.379 (0.011)
Table 8: Llama3.1-8B-Instruct on Formal Logic
  Round $t$   $p_{i,t-1}$   $\mathbb{E}[p_{i,t}\;|\;\mathcal{F}_{i,t-1}]$
  1   0.451   0.446 (0.016)
  2   0.443   0.452 (0.003)
  3   0.479   0.490 (0.016)
@@ Appendix G Closed-source LLM Evaluation
Appendix H Broader Impact, Limitations and Future Works
Broader Impact.
Limitations and Future Works.